OpenStudy (anonymous):

A half of half a number is 8. What is the number?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

32

jhonyy9 (jhonyy9):

32

OpenStudy (anonymous):

Pointless formalisation:\[x/4=8=>x=32\]

OpenStudy (anonymous):

x/2 = 8 multiplying both sides by 2... x = 16

OpenStudy (angela210793):

1/4X=8 so x=32

OpenStudy (anonymous):

Sorry misread the question ...................

OpenStudy (anonymous):

1/2 *1/2 * x = 1/4 *x = x/4 Equation ... x/4 = 8 multiply both sides by 4 x= 32